Calculate the de Broglie wavelength of a typical person walking through a doorway. Assume mass of...

  1. Calculate the de Broglie wavelength of a typical person walking through a doorway. Assume mass of the person to be 79 kg and her speed to be 1.4 m/s.
Answer #1

Given:

Mass of person ,m = 79 kg

Speed ,v = 1.4 m/s

de Broglie wavelength is given as:

= h/( m *v)

where h = Plank's constant = 6.626 * 10 -34 m​​​​​​2 kg/s

= 6.626 * 10 -34 / ( 79 * 1.4)

= 5.99095 * 10 -36​​​​ m